An EM-wave having speed 2•5x10^8units||Relative permeability of the medium is 0•8||Then find relative permittivity of the medium in which the wave is travelling?

Grade:10

1 Answers

Diali kundu
6 years ago
Velocity of em wave in vacuum is 3×10^8 m/s. Now in any other medium it is given by (3×10^8)/root (relative permeability of that medium ×relative permittivity of the medium )So putting the values we get,2.5×10^8=(3×10^8)/root (0.8×k)Solving we get k=1.8